# Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models

### v3

@@ Abstract
A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture not only stationary statistics but also responses to external perturbations. Current neural climate emulators seek to resolve the atmosphere–ocean system in all its complexity but often fail to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments. Response theory offers a rigorous framework to explore these limitations, moving beyond stationary statistics, probing causal mechanisms, and guiding model design. Using a simplified multiscale system, we argue that the skill of purely data-driven models in reproducing perturbed statistics depends critically on (i) identifying the proper variables to model, including spatial and temporal scales, and (ii) accurate stochastic parameterizations of unresolved processes. For studies of low-frequency climate dynamics, these insights motivate reduced-order neural models, tailored to specific processes, as valuable alternatives to general-purpose emulators. We illustrate this perspective with a reduced-order neural stochastic model designed to investigate responses of radiative fluxes to surface temperature perturbations. The model largely reproduces stationary statistics and enables the study of forced responses in both ensemble mean and variance, shifting the study of climate feedbacks from single trajectories to responses of probability distributions. These results underscore reduced-order models, augmented with modern neural networks and evaluated within the framework of response theory, as a practical strategy for causal inference and attribution studies in large-scale climate dynamics.
@@ I Introduction
A key challenge in current applied mathematics is to build data-driven models from partial observations of multiscale physical systems. Among the most challenging examples is the climate system: a forced, dissipative, and chaotic system whose dynamics span at least 15 orders of magnitude across spatial and temporal scales [1]. This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [2]. Seminal contributions include the work of Penland and collaborators [3, 4, 5], the multilevel modeling strategies of Kravtsov and colleagues [6, 7, 8, 9], and the contributions of Majda et al. [10, 11]. Recent proposal include Markov-chain models as in [12, 13, 14] and novel ideas from generative models [15]. Most prior work has focused on evaluating reduced-order models on equilibrium statistics and forecasting, with limited attention to responses to external perturbations, an exception being the contributions of Majda and collaborators (see e.g. [16]). Capturing such responses is key for building models that not only mimic the system’s dynamics but also faithfully capture causal relations across degrees of freedom [17, 18].
@@
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [19, 20, 21, 22, 23] and climate [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] systems. Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system. For example, [30] and [31] emulate the full ocean and atmosphere components of state-of-the-art climate models. Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons, offering a promising direction for efficient climate modeling. However, several limitations remain. Long rollouts suffer from error accumulation that can drive instabilities [34, 35]. Moreover, current emulators are unable to capture extreme events if not already present in the training set, preventing their use for rare-event analysis in a changing climate [36]. More relevant here, skill in reproducing stationary statistics does not guarantee accurate responses in perturbation experiments, restricting their use for causal studies of climate dynamics. This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related. The work of Van Loon et al. [37] illustrates this limitation in the context of the “pattern effect” problem [38, 39], i.e. the influence of sea surface temperature (SST) warming patterns on Earth’s energy imbalance. Using the ACE2 emulator [26] trained on ERA5 [40], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response. The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way. This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [30, 31, 41]. Even if such issues are addressed in specific cases, see [42] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Moreover, most current emulators represent only a subsystem (e.g., atmosphere or ocean), and limitations in responding to perturbations are expected to worsen in coupled models. Overall, limitations in capturing forced responses suggest that neural emulators are not yet ready for causal applications in climate dynamics.
@@
While the field of neural emulation is advancing rapidly, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings. In the first part of this paper, we build on a series of previous works by A. Vulpiani and collaborators on the limitations of data-driven modeling [43, 44, 45]. We consider a simplified stochastic dynamical system and ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations? In addressing this question, we focus on two relevant scenarios in data-driven-only modeling:
@@
  • Unknown equations, partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unknown. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
@@
To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [46, 47]. The model mimics the abstract structure of dynamical cores in large-scale climate models with a linear operator and an energy conserving nonlinear, quadratic operator. It provides a simplified yet relevant mathematical framework to unambiguously explore key limitations in capturing forced responses in climate emulators. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response to a general forcing can be computed by convolving the forcing with an impulse response operator [48, 49]. This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to small, general external perturbations [48], as well as the cause–effect relations among variables [17]. In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system. In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 50, 51], though typically without addressing responses to external perturbations. Therefore, we will argue that, in the case of low-frequency climate dynamics, a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
@@
We emphasize that the focus on a simplified dynamical system is not a limitation in this context. In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues of data-driven modeling that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings. As Smith (2002) [52] argues, while it is unreasonable to expect solutions to low-dimensional problems to generalize to high dimensions, so too it is unlikely that limitations observed in low-dimensional systems will vanish in high-dimensional, operational models. The first part of this paper is pedagogical in nature, aiming to bridge perspectives across researchers in statistical physics, climate science, and machine learning, while also providing practical methodological tools for the design and evaluation of data-driven models.
@@
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, when possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales. Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [50]. For an early instance of case (ii), see the work of Charney (1948) [53], Charney et al. (1950) [54] and the discussion in Dalmedico (2001) [51]. The example we consider is the “pattern effect” introduced above, which requires Green’s function experiments to quantify the response of top-of-atmosphere (TOA) radiative fluxes to perturbations in the sea surface temperature (SST) field. To this end, we develop a simplified stochastic neural model that emphasizes a large-scale representation of SST and global-mean net radiative flux, rather than attempting to resolve all spatial and temporal scales of variability as in current emulators. We discuss the coarse-graining and preprocessing choices relevant for model construction. The model largely reproduces both stationary and forced statistics and, importantly, enables Green’s function–like experiments to infer direct causal relations between SST and TOA radiative flux, a task where general-purpose emulators can perform poorly [37]. Crucially, we focus on responses in the ensemble mean and variance, rather than single trajectories as in most current studies, thereby extending the concept of climate feedback to include changes in distributional properties.
@@
The paper is organized as follows: in Section II we introduce the theoretical ideas underpinning this work. In Section III we discuss key conceptual difficulties in data-driven modeling of real-world dynamical systems. In Section IV we present a real-world application by focusing on radiative feedbacks to temperature perturbations. Conclusions then follow in Section V.
@@ II.1 Reduced-order stochastic climate models: theoretical and practical considerations
where $\mathbf{F}$, $\mathbf{L}$ and $\mathbf{n}(\mathbf{u})$ represent an external forcing, a linear and nonlinear operators, respectively. $\mathbf{u}$ is not necessarily finite dimensional, and in the context of coupled atmosphere-ocean variability, it could in principle represent the deviation of all fields, at all scales, from their climatology [4]. However, for obvious practical reasons, Eq. (1) needs to be simulated on finite spatial and temporal resolutions. In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is governed by a few dominant large-scale modes [4, 7, 48, 47, 55, 18]. The goal of reduced-order stochastic climate models is to develop low-dimensional models simulating large-scale, coarse-grained climate dynamics. The cumulative effect of unresolved fast variables, present at smaller spatial scales, is parametrized as stochastic terms. More formally, we decompose $\mathbf{u}$ into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, leading to an effective Langevin equation of the form:
@@
where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process. The noise covariance matrix $\mathbf{\Sigma}$ can be independent of $\mathbf{x}$, corresponding to additive noise, or state-dependent $\mathbf{\Sigma}(\mathbf{x})$ corresponding to multiplicative noise. In the latter case, the equation is here interpreted in the Itô sense. Equation (2) is mathematically justified in the limit of infinite time-scale separation between fast and slow modes [1]. In the most general (and often intractable) setting, the effective dynamics of the slow variables also involve memory effects induced by the fast degrees of freedom, leading to non-Markovian dynamics. Kondrashov et al. (2015) [7] addressed this problem in a practical and efficient way by introducing the framework of multilevel stochastic models. Nonetheless, appropriate coarse-graining procedures have been shown to effectively average out non-Markovian contributions in many climate applications, yielding useful Markovian approximations [3, 4]. We refer the reader to [2] for a review on the role of non-Markovian contributions in effective climate equations and to [56, 57] for specific examples.
Reduced-order stochastic climate models, as in Eq. (2), provide an efficient alternative to state-of-the-art climate models for studying low-frequency climate variability. They are much more computationally efficient and allow us for mechanistic understanding of climate dynamics by isolating core processes and examining their qualitative behavior. This has fueled the development of low-dimensional stochastic models learned either from data [58, 3, 4, 16, 10, 7, 2, 59] or from known equations [60, 61, 62, 63, 46], as well as inference-based techniques to infer interactions across modes of variability [64, 65, 66, 18].
@@
In spatiotemporal dynamical systems like climate, coarse-graining involves: (i) choosing a limited set of climate fields (e.g. temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales. Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained spatiotemporal fields obtained via (i–iii). The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [67, 8], by retaining the first $n<N$ of such bases. The correspondent principal components (PCs), serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as in [68, 18, 69] may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [70, 71]. Identifying the “right” coarse-grained representation remains one of the hardest challenges in data-driven modeling, as no general mathematical principle exists to guide such choices in high-dimensional dynamical systems [43, 17, 17].
@@
Given the coarse-grained data $\mathbf{x}$, the goal is to infer a stochastic Markovian model as in Eq. (2). We summarize several traditional approaches that motivate our work, while acknowledging that many other methods exist beyond those discussed here. The simplest example is the Linear Inverse Model (LIM) proposed by Penland [3], where the drift is linear and the noise is additive, both estimated directly from data. LIMs have shown forecasting skill for sea surface temperature anomalies comparable to climate models [4, 5, 72, 73], and extensions exist to handle colored noise [74]. Kravtsov et al. (2005) and Kondrashov et al. (2015) [6, 7] generalized this framework through multilevel regression, incorporating quadratic drift terms and non-Markovian effects. These models perform well for stationary statistics [8] and forecasting [75]. The work of Majda and Harlim extended the multilevel regression framework by proposing stability constraints [11]. A theoretically grounded alternative is the approach of Majda, Timofeyev, and Vanden-Eijnden [60, 61, 62, 63], dubbed MTV by the authors, which derives effective equations for the coarse-grained variables $\mathbf{x}$ rather than assuming them. MTV has been successful for simplified models, though less so for complex climate systems, likely due to choices of coarse-graining steps (i-iii) shown above rather than intrinsic limitations, see also discussion in [8]. We will expand on this approach in Section III and Appendix B.
Overall, these works demonstrated that robust reduced-order stochastic climate models can be built through relatively simple regression strategies. In this paper, we build on these insights and investigate how simple neural architectures can infer both the drift and state-dependent noise in Eq. (2), with minimal assumptions. Our neural models retain a decomposition into linear and nonlinear terms, $\mathbf{L}$ and $\mathbf{n}(\mathbf{x})$, as in the general representation of dynamical systems in Eq. (1). Furthermore, we will propose simple solutions to fit multiplicative (state-dependent) noise in high dimensions.

### v4

# Probing forced responses and causality in data-driven climate emulators: conceptual limitations and the role of reduced-order models

@@ Abstract
A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture both stationary statistics and responses to external perturbations. Current neural climate emulators aim to resolve the atmosphere–ocean system in all its complexity but often fail to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments. To investigate the origin of these limitations, we first focus on a simplified dynamical system that retains key features of climate variability. We interpret the results through linear response theory, providing a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms. We argue that the ability of multiscale systems’ emulators to reproduce perturbed statistics depends critically on (i) identifying an appropriate coarse-grained representation and (ii) careful parameterizations of unresolved processes. For low-frequency climate dynamics, these insights highlight reduced-order models, tailored to specific processes and scales, as valuable alternatives to general-purpose emulators. We next consider a real-world application, developing a neural model to investigate the joint variability of the surface temperature field and radiative fluxes. The model infers a multiplicative noise process directly from data, largely reproduces the system’s probability distribution, and enables causal studies through forced responses. We discuss its limitations and outline directions for future work. These results expose fundamental challenges in data-driven modeling of multiscale physical systems and underscore the value of coarse-grained, stochastic approaches, with response theory as a principled framework to guide model design.
@@ I Introduction
A key challenge in current applied mathematics is to build data-driven models from partial observations of multiscale physical systems. One of the most complex examples is the climate system: a forced, dissipative, and chaotic system whose dynamics spans at least 15 orders of magnitude across spatial and temporal scales [GhilLucarini]. This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [LucariniChekroun]. Seminal contributions include the work of Penland and collaborators [Penland89, Penland95, PENLAND1996534], the strategy of Kravtsov and colleagues [KRAVTSOV, KONDRASHOV201533, STROUNINE2010145, Agarwal], and the contributions of Majda et al. [majdaVsGhil, majdaConstraints]. Recent proposals include Markov-chain and generative approaches [Andre1, Andre2, GIORGINI2024134393, GiorginiScore, CycloLudoAndre], as well as autoencoder-based, latent-space formulations [PierreLatent]. Most prior work has focused on evaluating reduced-order models on equilibrium statistics and forecasting, with limited attention to responses to external perturbations. Notable exceptions include the contributions of Majda and Qi [majdaSIAM, QiMajdaChaos]. Capturing such responses is key for building models that not only mimic the system’s dynamics but also faithfully capture causal relations across degrees of freedom [Baldovin, FabCausal].
@@
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [Pathak2022, gencast, Lam2023, Bi2023, fourcastnet3] and climate [Kochkov2023, Watt-Meyer2023, Watt-Meyer2024, Lucie, OceanNet, Bire, Samudra, Camulator, lucie3D, SamudraACE]. Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system. For example, [Samudra] and [Camulator] emulate the full ocean and atmosphere components of state-of-the-art climate models. Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons. They can potentially be used for tuning and sensitivity analysis [M2Lines], and to improve the study of rare events [Lancelin], offering a promising direction for fast and efficient tools to study climate. However, several limitations remain: for example, long rollouts can suffer from error accumulation that can lead to instabilities [Chris, PedramStability]; rare events are not captured if absent from the training set, limiting their applicability for rare-event analysis in a changing climate [PedramPNAS]; key atmospheric modes, such as the Quasi-Biennial Oscillation and the Southern Annular Mode, are not well represented [Baxter]. A limitation particularly relevant for climate science, as opposed to weather forecasting, is that high skill in reproducing stationary statistics does not ensure accurate responses in perturbation experiments, thereby limiting the use of such models for causal studies of climate dynamics. This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related. The work of Van Loon et al. [Senne] illustrates this limitation in the context of the “pattern effect” problem [blochjohnsonetal2024, Dong2019AttributingPacific], i.e. the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance. Using the ACE2 emulator [Watt-Meyer2024] trained on the ERA5 reanalysis [era5], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response. The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way. This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [Samudra, Camulator, ZhangACE]. Even if such issues are addressed in specific cases, as shown by Wu et al (2025) [Wu] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Most current climate emulators represent only a subsystem (e.g. atmosphere or ocean) and limitations in capturing forced responses are likely to become more pronounced in coupled settings. Additionally, emulators’ skill in capturing forced responses is often assessed using single trajectories. In climate applications, however, the relevant quantities are probability distributions, particularly the responses of the mean and variance. We will show that focusing on the skill in capturing variance responses can reveal important limitations otheriwse hidden when examining only the mean response. Taken together, these issues indicate that neural emulators may not yet be reliable general-purpose tools for causal studies of climate dynamics.
@@
While the field of neural emulation is advancing rapidly and yielding impressive results, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings. In the first part of this paper, we build on a series of previous works by Vulpiani and collaborators on the limitations of data-driven modeling [FutureFromPast, Hosni, BaldovinEntropy]. We consider a simplified stochastic dynamical system and ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture responses to general external perturbations? In addressing this question, we focus on two relevant scenarios in data-driven-only modeling:
@@
  • Unknown equations, partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unobserved [e20070509], or even unknown [BaldovinEntropy]. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
@@
To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [NormalForms, Majda2010]. The model mimics the abstract structure of dynamical cores of large-scale climate models with a linear operator and energy conserving quadratic nonlinearities. It provides a simplified yet relevant mathematical framework to unambiguously explore fundamental (rather than technical) limitations in capturing forced responses in climate emulators. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response can be computed by convolving the forcing with an impulse response operator [MajdaBook, Marconi]. This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to general external perturbations. Therefore, linear response theory is here proposed as a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms [Baldovin]. In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system. In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [PENLAND1996534, Gammaitoni, Dalmedico], though typically without addressing responses to external perturbations. Therefore, we will argue that, in the case of low-frequency climate dynamics, a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
@@
We emphasize that the focus on a simplified dynamical system is not a limitation in this context. In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues of data-driven modeling that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings [Lenny2002]. The first part of this paper is pedagogical in nature, aiming to bridge perspectives across researchers in statistical physics, climate science, and machine learning, while also providing practical methodological tools for the design and evaluation of data-driven models.
@@
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, when possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) when even in the presence of all variables the system encompasses a huge number of spatial and temporal scales. Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [Gammaitoni]. For an early instance of case (ii), see the work of Charney (1948) [charney1949], Charney et al. (1950) [neumann] and the discussion in Dalmedico (2001) [Dalmedico]. The example we consider is the “pattern effect” introduced above, which requires Green’s function experiments to quantify the response of top-of-atmosphere (TOA) radiative fluxes to perturbations in the sea surface temperature (SST) field. To this end, we develop a simplified stochastic neural model that emphasizes a large-scale representation of SST and global-mean net radiative flux, rather than attempting to resolve all spatial and temporal scales of variability as in current emulators. We discuss the coarse-graining choices relevant for model construction, as well as a practical and efficient way to fit multiplicative noise from data. The model largely reproduces stationary statistics, especially the most energetic SST mode, corresponding to the El Niño Southern Oscillation [ENSOcomplexity], and it enables Green’s function–like experiments to infer direct causal relations between SST and TOA radiative flux. The discussion also addresses the model’s limitations and potential methodological developments.
@@
The paper is organized as follows: in Section II we introduce the theoretical ideas underpinning this work. In Section III we discuss key conceptual limitations in data-driven modeling of real-world dynamical systems. In Section IV we present a real-world application. A summary of the main findings and proposals follow in Section V and Section VI.
@@ II.1 Reduced-order stochastic climate models: theoretical and practical considerations
where $\mathbf{F}$ is an external forcing, and $\mathbf{L}$, $\mathbf{n}(\mathbf{u})$ represent linear and nonlinear operators, respectively. $\mathbf{u}$ is not necessarily finite dimensional, and in the context of coupled atmosphere-ocean variability, it could in principle represent the deviation of all fields, at all scales, from their climatology [Penland95]. However, for obvious practical reasons, Eq. (1) needs to be simulated on finite spatial and temporal resolutions. In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is governed by a few dominant large-scale modes [Penland95, KONDRASHOV201533, MajdaBook, Majda2010, Dubrulle, FabCausal]. The goal of reduced-order stochastic climate models is to capture the evolution of these modes and their interactions. This modeling choice allows us to move away from a grid-based representation of the system, emphasizing resolution-independent modes as fundamental components of the system. The cumulative effect of unresolved fast variables, present at smaller spatial scales, is parametrized as stochastic terms. More formally, we decompose $\mathbf{u}$ into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, leading to an effective Langevin equation of the form:
@@
where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process. The noise covariance matrix $\mathbf{\Sigma}$ can be independent of $\mathbf{x}$, corresponding to additive noise, or state-dependent $\mathbf{\Sigma}(\mathbf{x})$ corresponding to multiplicative noise. In the latter case, the equation is here interpreted in the Itô sense. Equation (2) is mathematically justified in the limit of infinite time-scale separation between fast and slow modes [GhilLucarini]. In the most general (and often intractable) setting, the effective dynamics of the slow variables also involve memory effects induced by the fast degrees of freedom, leading to non-Markovian dynamics. Kondrashov et al. (2015) [KONDRASHOV201533] addressed this problem in a practical and efficient way by introducing the framework of multilevel stochastic models. Nonetheless, appropriate coarse-graining procedures have been shown to effectively average out non-Markovian contributions in many climate applications, yielding useful Markovian approximations [Penland89, Penland95]. We refer the reader to [LucariniChekroun] for a review on the role of non-Markovian contributions in effective climate equations and to [Wouters, Vissio2018] for specific examples.
Reduced-order stochastic climate models, as in Eq. (2), provide an efficient alternative to state-of-the-art climate models for studying low-frequency climate variability. They are much more computationally efficient and allow us for mechanistic understanding of climate dynamics by isolating core processes and examining their qualitative behavior. This helps bridging the gap between simulations and understanding [HeldGap]. This has fueled the development of low-dimensional stochastic models learned either from data [Hasselmann, Penland89, Penland95, majdaSIAM, majdaVsGhil, KONDRASHOV201533, LucariniChekroun, nanChen] or from known equations [MTV1, MTV2, MTV3, MTV4, NormalForms], as well as inference-based techniques to infer interactions across modes of variability [Kane, Donges, Ilias_1, FabCausal].
@@
In spatiotemporal dynamical systems like climate, coarse-graining involves: (i) choosing a limited set of climate fields (e.g. temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales. Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained representation obtained via (i–iii). The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [MajdaPerspective, STROUNINE2010145], by retaining the first $n<N$ of such bases. The correspondent principal components (PCs), serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as principal interaction patterns [CROMMELIN], community detection [FabCausal] and autoencoders [latentPierre, PierreLatent] may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [HeldFDT, Pedram2]. Identifying the “right” coarse-grained representation depends on the specific goals of the researcher and remains one of the most difficult challenges in data-driven modeling, as no general mathematical principle exists to guide such choices from data alone in high-dimensional dynamical systems [FutureFromPast, Baldovin, Baldovin].
@@
Given the coarse-grained data $\mathbf{x}$, the goal is to infer a stochastic Markovian model as in Eq. (2). We summarize several traditional approaches that motivated our work, while acknowledging that many other methods exist beyond those discussed here. The simplest example is the Linear Inverse Model (LIM) proposed by Penland [Penland89], where the drift is linear and the noise is additive, both estimated directly from data. LIMs have shown forecasting skill for sea surface temperature anomalies comparable to climate models [Penland95, PENLAND1996534, Penland_Fowler, Capotondi]. Recent developments extended the LIM framework to include colored noise [coloredLIM] and nonlinearities [LIM_nonlinear]. Kravtsov et al. (2005) and Kondrashov et al. (2015) [KRAVTSOV, KONDRASHOV201533] generalized this framework through multilevel regression, incorporating quadratic drift terms and non-Markovian effects. These models perform well for stationary statistics [STROUNINE2010145] and forecasting [KONDRAENSO]. The work of Majda and Harlim extended the multilevel regression framework by proposing stability constraints [majdaConstraints]. A theoretically grounded alternative is the approach of Majda, Timofeyev, and Vanden-Eijnden [MTV1, MTV2, MTV3, MTV4], dubbed MTV by the authors, which derives effective equations for the coarse-grained variables $\mathbf{x}$ rather than assuming them. MTV has been successful for simplified models, though less so for complex climate systems, likely due to choices of coarse-graining steps (i-iii) shown above rather than intrinsic limitations, see also discussion in [STROUNINE2010145]. We will expand on this approach in Section III and Appendix B.
Overall, these works demonstrated that robust reduced-order stochastic climate models can be built through relatively simple regression strategies. In this paper, we build on these insights and investigate how simple neural architectures can infer both the drift and state-dependent noise in Eq. (2) with minimal assumptions. Our neural models retain a decomposition into linear and nonlinear terms, $\mathbf{L}$ and $\mathbf{n}(\mathbf{x})$, as in the general representation of dynamical systems in Eq. (1).